OpenStudy (cherrilyn):

evaluate the integral by completing the square and using trig sub:

OpenStudy (anonymous):

The function?

OpenStudy (cherrilyn):

\[\int\limits_{?}^{?}\sqrt{(x ^{2}-4x+7)dx}\] and \[\int\limits_{}^{?} dx/(x ^{2} +6x + 6)^{2}\]

OpenStudy (cherrilyn):

for the first one I know the completed square is (x^2-4x+4)-4+7 so I get (x-2)^2+3

OpenStudy (cherrilyn):

then I go all the way down to u sub which leads me to \[\int\limits_{}^{}\sqrt{u ^{2}+3}du\]

OpenStudy (anonymous):

You would get better help if you pose one quetion at a time. Or else results become difficult to follow. So which one are you working on?

OpenStudy (cherrilyn):

the first one

OpenStudy (anonymous):

Only place I have ever seen this process address is in this MIT lecture beginning at about 37:52 http://www.youtube.com/watch?v=CXKoCMVqM9s&feature=BFa&list=PL590CCC2BC5AF3BC1&index=25

OpenStudy (cherrilyn):

thanks!! :D